What can be said about an endothermic reaction with a negative entropy change?

The reaction is

What can be said about an endothermic reaction with a negative entropy change?

spontaneous at all temperatures.
spontaneous at high temperatures.
spontaneous at low temperatures.
spontaneous in the reverse direction at all temperatures.
nonspontaneous in either direction at all temperatures.
Answer #1

Ans: spontaneous in the reverse direction at all temperatures.

?G = ?H -T?S

For spontaneous reactions, ?G = -ve

Given that endothermic reaction with a negative entropy change.

?H = +ve ?S= -ve

?G = ?H -T?S = +ve - T (-ve) = +ve

Therefore, the given reaction is non-spontaneous in forward direction.

Reverse direction:

Given that endothermic reaction with a negative entropy change. Then, ?H = +ve ?S= -ve

For reverse reaction ?H = -ve ?S= +ve

?G = ?H -T?S = -ve - T (+ve) = -ve

?G = -ve

Therefore, the given reaction is spontaneous in reverse direction.